UT 13 Sep 2021




Night Staff:

OSA: D.G. Huerta

ISp: Mtn Staff

ISA/Observers: J.Power

Technical Wizard: J. Hill (Tucson)

Mtn Manager: M. Wagner

Sunset: 18:34

Twlight Ends: 19:52

Twilght Starts: 04:39

Sunrise: 05:57

Starting Config: LUCI-Bino (LUCI-1 authorized)

Night Log:


Proposed Plan of Action:

  • LUCI-1:
    • LUCI-1 Dual BLR target - twilight until 30deg limit (Rothberg) - until 20:45 local time
      • Take Telluric first as soon as AGW-1 can wave-front sense
  • Reconfigure for MODS-Bino
    • DONE Rho Oph - observations
    • DONE Spectrophotometric Standard
  • If Transforms ready proceed to appropriate instrument to collect Transform Confirmatory data - (whichever is ready to go) - highest priority
  • MODS-Bino
    • DONE Update Transforms in agw.ini for both MODS
    • DONE Update LZP for MODS
      • Make a backup of PCSInstrument.conf
      • Copy to PCSInstrument.conf
    • DONE Restart systems (GCS, PCS) and Restart MODS services (restart of PCS only)
    • DONE Confirmatory AGW Transform - 1.5hr - Highest Priority seeing 1.2" or better
      • Fields:
        • StoneO 21 37 31 +00 31 55 - airmass 1.3 at 20:45
        • DONE StoneP 22 35 49 +01 31 33
        • StoneA 00 54 02 -00 32 34
      • DONE Re-run MODS MOS Astrometric masks to check the Rot ZP
    • DONE Dither test MODS (1hr), scripts in /home/lbto/Service_Observing/Restart_2021/MODS/DitherTest
    • Spectrophotometric stars (if clear any seeing)
    • Standard Stars (if clear any seeing)
    • Pointing Model (poor weather and poor seeing)
    • NS Longslit - low priority
    • Photometric and Spectro-photometric standard (if clear and poor seeing)
    • Spectrophotometric standard all modes
      • Grating: Dual, Blue, Red
      • Prism: Dual, Blue, Red
  • LUCI:
    • AGW-1/2 Transform Confirmatory Data - 1.5hr - Highest Priority
      • Fields:
      • StoneO 21 37 31 +00 31 55 Stone_O_m12.list - airmass 1.3 at 20:45
      • StoneP 22 35 49 +01 31 33 Stone_P1.list
      • StoneA 00 54 02 -00 32 34 Stone_A3.list
    • Photometric Standards (clear any seeing)
  • Mixed Mode Test - if time
    • LBC-Red & MODS-1
      • Test whether LBC can send a 2nd script (Dec -90) and everything works well for both sides
  • LBC:
    • Photometric Standards (clear any seeing)
    • Spectrophotometric stars (clear any seeing)
    • Pointing Model LBC (poor seeing, clouds) - 2-3hrs
      • Turn on PCS logging
      • Turn on TMS passive mode
      • Need to collimate at each star - or - collimate every few stars
      • Chip 2 only exposure
      • File Location:
        • /home/lbto/Calib_OBs/PtStars_10m_OBs/VV
        • PtStarrs: /home/lbto/POINTING/PtStars_10m/PtStars_10m.dat
    • If conditions too poor for other stuff - Pointing Model in Active mode too
    • General TMS Active mode testing (any conditions)
      • run any Science-like OB in the LBC Restart OT file
      • Run for 1-1.5 hours
  • Poor Weather (if pointing models are exhausted)
    • LUCI-1 local mergers
    • Service_Observing/Restart_2021/LUCI/PoorWeather/COBandhead_Mergers_V17.xml - OT file
    • finder charts in sub-directory (but they are so large you can't miss 'em)

Current Status

  • LBCs
    • LBC-Blue - available
    • LBC-Red - available
  • MODS
    • MODS-1 - issues with M1RC headers & filenumbers short 1 digit
      • AGW-5 - available
    • MODS-2 - available
      • AGW-6 - available
  • LUCI-1 - available
    • AGW-1 available
  • LUCI-2 offline
    • AGW-2 available
  • PEPSI - available
    • AGW-7/8 available


Closed Dome Activities

  • MODS-1 Full set of Cals
    • Biases
      • 1x1
        • 8k - COMPLETED
      • 1x2
      • 2x2
    • Grating
      • Dual
        • 0.8" slit, 0.3" slit, slitless pixel flats, arcs - COMPLETED
      • Red
        • 0.8" slit, slitless pixel flats, arcs - COMPLETED
      • Blue
    • Imaging
      • Dual
      • Red
      • Blue
    • Prism
      • Dual
      • Red
      • Blue
  • MODS-2 Cals
    • Bias
      • 1x2
      • 2x2
    • Prism
      • Dual
      • Red
      • Blue
  • LUCI-1 Cals
    • Imaging N375 Flats
      • Narrow, medium-band filters
      • Broad zJHKKs and HeI completed
    • N1.8 G210 CWL=2.3 K-band 0.5" cals
    • N1.8 G200 CWL=1.93 HKspec 0.5" cals
  • ARGOS Cal-Unit test for M3 alignment and pupil wobble checkout (alternate to on-sky method)
  • LUCI-1 Only Calibrations
    • Flexure data for G150
      • /home/dthomspons/Luci/FlexureCompensation
      • README - Calib15 - 2hrs
      • Ks filter, Argon lamp, optical sieve mask
      • take dark frame
      • Manual Control MCSGUI for elevations
    • Wavelength data for G150
      • G150 Ks, Argon, take image through slit, then put in grating, line should show up where the slit is located (within a few pixels of the image of the slit)
      • Check a few lines (2 or 3)

Night Summary.

For the night of: Sun, 2021 Sep 12 ---> Mon, 2021 Sep 13
Local midnight = 2021 Sep 13,  7 hr UT, or JD 2459470.792
Local Mean Sidereal Time at midnight =  23 10 25.7 

Sunset (  500 m horizon):   18 33 MST; Sunrise:   5 59 MST
Evening twilight:  19 52 MST;  LMST at evening twilight:  19 02
Morning twilight:   4 39 MST;  LMST at morning twilight:   3 50
12-degr twilight: 19 23 MST -->  5 08 MST; night center:   0 16 MST 

Moonset :  22 33 MST   
Moon at civil midnight: illuminated fraction 0.428
0.6 days before first quarter, RA and dec:  16 43 21, -23 34.9 
The sun is down for 11.4 hr;  8.8 hr from eve->morn 18 deg twilight.
 6.1 dark hours after moonset and before beginning of twilight.

MODS, LUCI1, and LBS strated and checked out. Running MODS bias all, LBC's 25 bias, and LUCI darks.

1:30 RPC error at 16/25 for red images. lbckill/lbc start. Banding on blue. Running a 2 bias to ensure all is running well since we are close to sunset, and success.

1:39 Particles are riding the limits for opening, triggering the alarm handler. David is chacking conditions and there is a layer of dust. We will be holding off.

1:46 Since we are holding off on opening I will rerun my 25 bias script for LBC's. Low level banding still present on LBCB

3:07 Particles continue to spike over the limits. We are running too low on the for the Dual BLR target. Reconfiguring to binocular MODS.

4:04 David is opening. Particles have dropped below the limit. The moon is just up on the western hoorizon at 0.428% illumniation. We are still going to try for RS Oph first but it will have moon light.

RhoOph

4:28 David is pointed and Collimated. Seeing is 0.9-1.1" on the guiders. David had some weirdness related to his gnome session. Sending preset to RS Oph with PA 44.

Target is quite saturated, even in 0.5 sec acquisition. MODS 1 alignment is not ideal but too saturated and object bouces with guider from imge to image more than anything.
Spectrum saturated in red emission line. Reducing exposure time from 2 seconds to 0.5.

Looks like 0.5sec is ok

mods1r.20210913.0028-34
mods2r.20210913.0024-30
mods[1/2]b.20210913.0021-27

5:08 Sending preset to Feige110

mods1r.20210913.0037-39
mods2r.20210913.0033-35
mods[1/2]b.20210913.0028-30

MODS Transform Verfication

5:37 Slewing to Stone_P1 for transform data collection. Taking a quick R image before we start to configure mods for dual red only imaging, r, 5 sec.

  1. Update PCSInstrument.conf and bounce PCS
  2. stop mods agw services
  3. update agw.ini
  4. restart MODS agw services
  5. David is sending acquire preset and refining pointing after restart of PCS.
  6. Transform list:Stone_P1_mods.list. Made the list for mods patrol field at angle 245,255, 265, StoneP 22 35 49 +01 31 33,
    • Seeing 0.9-1" on the guider
    • I didn't realize the DX would take blue imaging and not red. I had configured for red imaging. Restarting set up for blue only imaging with g.
    • 6:24 Seeing improving, 0.75-0.85", mostly clear
    • GS#4 appears to be a double. Not a good source
    • 6:41 and seeing is back up to 1". variable.
    • data is located in /home/lbto/data/20210913/ , obtained 16 stars for SX and 14 with DX

RotZP verification

7:07 Sending preset to StoneA for Astrometric mask acquisition to verify rotZP. Seeing 0.7" on the guider. Mostly clear. Mask mods.37.StoneA.mms
MODS1:

Computed Mask Alignment Offset:
dX = 0.201 arcsec
dY = 0.734 arcsec
dPA = -0.046 deg

MODS1 Offset Command:
offsetpointing -0.0457 0.201 0.734 detxy rel

MODS2:

Computed Mask Alignment Offset:
dX = 4.317 arcsec
dY = -0.924 arcsec
dPA = -0.006 deg

MODS2 Offset Command:
offsetpointing -0.0061 4.317 -0.924 detxy rel

RotZP looks good. Moving to dither test

MODS Dither Test

7:30 Slew to M52 for Dither test. Seeing has blown right up to 2"

There is not finder chart, so i'm not sure which 2 stars I'm to be placing 5" from the bottom. Shifted absolut Y -6" which places 2 stars roughly 5" from the bottom edge of the pointing mask.

Problem with the dither script? Looks ok. Keeps dumping down and trying to start next exposuring while ex[posuing. and offseting during exposures. Abort and restart. WFS images look chopped off on the top. Second try was fine. Perhaps manually sedning offsets and taking exposures I created a conflict.

8:07 Collimation is really running away on the right side. aborting and David is removing the corrections on the primary. That appears to have fixed it and we have what looks like a star again rather than a jabba the hut.

8:20 DX is again looking really bad. David removed corrections and it looks worse. Estimated seeing is 1.25" on DX and 0.85" on SX. Mirrors are in equillibrium and david reports no large temperature variations that would explain this.

8:32 It looks like condensation is building on the MODS2 images. Collimation on DX is stuggling and

Blob forming during dither test ~ 8:30 UT on DX:
MODS_BLOB_DURING_DITHER_TEST.png

Humidity 40.9%

The shackhartmann on both GCS images are cut off:
Shack Hartman Vignette.png

David has had to regularly remove the active optics on the primary regularly to maintain collimation.

The blob appears to be shifting as we expose
  • Blob on DX appears to be shifting 8:49 UT:
    Blob2 Dither Test.png

LBC Pointing Model

8:53 Moving up to zenith and reconfiguring to LBC's. Andrew has powered on lasers for Passive TMS.

8:26 LBC's authorized. Starting pointing log and starting pointing model.

Want to start with somethign a little higher for first collimation of the night.

Slewing to star 17, el 52, az 29, Runaway on red. David clearing optics.

8:49 Particle monitor is over the limit. We are closing. Stopped pointing log (we never really started).

Looking into the Blob

The blob first appears as a donut in mods2 images 0058, visible on the pointing mask.
mage first appeared on image 58, visible on pointing mask.:
First appearance of the blob on image 0058.png

The images with TELAZ TELALT ROTANGLE including the one prior are:

mods2r/b.20210913.0057.fits -15.37836 59.06616 -1.10065 M52 dither test
mods2r/b.20210913.0058.fits -15.66230 58.98777 -1.65368 M52 dither test
mods2r/b.20210913.0059.fits -15.92712 58.91301 -2.17011 M52 dither test
mods2r/b.20210913.0060.fits -16.19692 58.83490 -2.69919 M52 dither test
mods2r/b.20210913.0061.fits -16.46938 58.75417 -3.23461 M52 dither test
mods2r/b.20210913.0062.fits -16.72941 58.67540 -3.74799 M52 dither test
mods2r/b.20210913.0063.fits -16.98701 58.59547 -4.25876 M52 dither test
mods2r/b.20210913.0064.fits -17.24218 58.51505 -4.76561 M52 dither test
mods2r/b.20210913.0065.fits -17.49872 58.43187 -5.27803 M52 dither test
mods2r/b.20210913.0066.fits -17.74730 58.34926 -5.77659 M52 dither test
mods2r/b.20210913.0067.fits -17.99635 58.26494 -6.27877 M52 dither test
mods2r/b.20210913.0068.fits -18.26056 58.17346 -6.81283 M52 dither test
mods2r/b.20210913.0069.fits -18.51292 58.08394 -7.32671 M52 dither test
mods2r/b.20210913.0070.fits -18.75712 57.99554 -7.82465 M52 dither test
mods2r/b.20210913.0071.fits -18.99701 57.90688 -8.31709 M52 dither test
mods2r/b.20210913.0072.fits -19.23684 57.81614 -8.81232 M52 dither test
mods2r/b.20210913.0073.fits -19.47531 57.72438 -9.30507 M52 dither test
mods2r/b.20210913.0074.fits -19.71071 57.63148 -9.79619 M52 dither test
mods2r/b.20210913.0075.fits -19.71071 57.63148 -9.79619 M52 dither test
mods2r/b.20210913.0076.fits -20.22946 57.42060 -10.88414 M52 dither test
mods2r/b.20210913.0077.fits -20.45577 57.32544 -11.36291 M52 dither test
mods2r/b.20210913.0078.fits -20.68289 57.22837 -11.84527 M52 dither test
mods2r/b.20210913.0079.fits -20.91608 57.12632 -12.34287 M52 dither test
mods2r/b.20210913.0080.fits -21.13872 57.02708 -12.82121 M52 dither test

Could it have been a bug taking a stroll?

Taking mods imaging flats. Well, my bug theory may hold weight. It looks like something decided to walk all over the MODS2 optics
Something has made a mess of things - mods2 g and r imaging flats:
ImagingFlats.png

Taking a blue only imaging glat to rule out the shmutz is on the dichroic

Phot Standard

11:50 Particles have dropped. Not much time remaining. We will try and sneak in a pho stnandard but likely will just be open to get sky flats.

11:58 slewing to J2341399+453900 and running dohybrid.

11:06 its already 12 degree. I'm going to a blank field for sky flats.

LBC SkyFlats

12:12 slewing to blank 04+25. Primary Mirror Panic'd

12:26 Primary recovered. Sending preset to field for Flats.

Obtained a good set at PA0 and PA180 of LBCB r & LBCR V

Closed up, david reconfigured with LBC's out of the beam.

Taking 25 LBC biases

Think I found the source of the weirdness with the script. I had an unclosed window from a script run earlier hiding behind one of my ds9 windows. Rick mentioned those can cause oddities.

Transform and ZP for MODS left in place

-- BarryRothberg - 12 Sep 2021

  • I

*
I Attachment Action Size Date Who Comment
Blob2 Dither Test.pngpng Blob2 Dither Test.png manage 676 K 13 Sep 2021 - 08:49 JenniferPower Blob on DX appears to be shifting 8:49 UT
First appearance of the blob on image 0058.pngpng First appearance of the blob on image 0058.png manage 281 K 13 Sep 2021 - 10:17 JenniferPower Image first appeared on image 58, visible on pointing mask.
ImagingFlats.pngpng ImagingFlats.png manage 955 K 13 Sep 2021 - 10:52 JenniferPower Something has made a mess of things - mods2 g and r imaging flats
MODS_BLOB_DURING_DITHER_TEST.pngpng MODS_BLOB_DURING_DITHER_TEST.png manage 674 K 13 Sep 2021 - 08:37 JenniferPower Blob forming during dither test ~ 8:30 UT on DX
Shack Hartman Vignette.pngpng Shack Hartman Vignette.png manage 635 K 13 Sep 2021 - 08:40 JenniferPower The shackhartmann on both GCS images are cut off
Topic revision: r9 - 16 Sep 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback